# Beyond twisted maps: applications to motivic integration

### v2

@@ 1. Introduction
If $Y$ is a log-terminal variety with quotient singularities, its stringy Hodge numbers have a beautiful interpretation as orbifold Hodge numbers of a smooth Deligne–Mumford stack $\mathcal{X}$ where $\mathcal{X}\to Y$ is a canonical crepant resolution. There are two major obstructions to finding a similar interpretation for the stringy Hodge numbers when $Y$ has worse than quotient singularities. The first is that $Y$ rarely admits a crepant resolution by a Deligne–Mumford stack. In [SU23a], we were able to overcome this obstruction by constructing crepant resolutions for all log-terminal varieties using Artin stacks. This leads to the second major obstruction: Artin stacks (other than Deligne–Mumford stacks) are not separated, creating major difficulties in applying previous techniques in motivic integration. In this paper we introduce a versatile framework that overcomes this second obstruction.
@@ 1.1. Overview
While crepant resolutions by varieties or Deligne–Mumford stacks are rare, we proved in recent work [SU23a] that crepant resolutions by Artin stacks hold in great generality: every log-terminal variety $Y$ admits a crepant resolution by a smooth Artin stack $\mathcal{X}$. Since log-terminal singularities are precisely the class for which Batyrev defined stringy Hodge numbers [Bat98], this shows that resolutions by smooth Artin stacks provide a natural setting by which to study stringy invariants. Furthermore, log-terminal singularities are a natural class arising in the context of the minimal model program (MMP). Since motivic integration and the MMP are largely controlled by maps from discs and curves, for applications to birational geometry and motivic McKay correspondences, it is particularly important to understand how maps from curves or arcs on $Y$ lift to twisted versions on $\mathcal{X}$. One is therefore in need of new tools for lifting curves and arcs via resolutions by Artin stacks.
@@
Furthermore, we provide a versatile framework for working with warped maps. Indeed, while the valuative criterion and the lack of properness for good moduli spaces provides the impetus for the construction of warped maps, the most prominent feature of our paper is the construction of an Artin stack $\mathscr{W}\!(\mathcal{X})$ which parameterizes warped maps, i.e., a warped map from $T$ to $\mathcal{X}$ is given by a usual map $T\to\mathscr{W}\!(\mathcal{X})$, see Theorem 1.1. Thus, this allows one to bootstrap results from usual maps to warped maps. This bootstrapping technique is essential in our main applications to motivic integration and the McKay correspondence (Theorems 1.4 and 1.5), and is analogous to Olsson’s work in logarithmic geometry [Ols03], where he showed that logarithmic morphisms $S\to Z$ are the same as usual morphisms $S\to\mathcal{L}og_{Z}$ to an auxiliary stack, and hence properties of log maps are inherited from properties of usual maps to stacks.
@@
With the theory of warped maps in hand, we prove a motivic change of variables formula for Artin stacks (see Theorem 1.4) and give a McKay correspondence for linearly reductive groups (see Theorem 1.5). Our change of variables formula for Artin stacks applies, in particular, to crepant resolutions of all log-terminal singularities, as constructed in our previous work [SU23a]. While other motivic change of variables formulas for stacks have appeared in the literature, they are all either in the Deligne–Mumford setting or involve making choices, e.g., choosing lifts to arcs of $\mathcal{X}$. Such choices do not always exist and even when they do, there can be infinitely many such choices. On the other hand, our motivic change of variables formula here is completely canonical: by our valuative criterion (Theorem 1.2) outside of a set of measure zero, any arc $\varphi$ of $Y$ lifts canonically to a warped arc $\varphi^{\operatorname{can}}$ of $\mathcal{X}$. Our motivic change of variables formula then expresses any motivic integral over arcs $\varphi$ of $Y$ as an explicit motivic integral over the warped arcs $\varphi^{\operatorname{can}}$ of $\mathcal{X}$.
@@
We wish to emphasize that in addition to our motivic change of variables formula being a novel result, the proof technique is unlike any that previously existed, even when one restricts to the case of Deligne–Mumford stacks. Indeed, for Yasuda to construct a motivic measure for twisted arcs, he needed to first construct spaces parametrizing twisted jets and arcs and then get enough explicit control of these spaces in order to construct the measure. In comparison, our main theorem (Theorem 1.1) showing $\mathscr{W}\!(\mathcal{X})$ is an Artin stack allows us to completely sidestep the need to construct a motivic measure for warped arcs: since a warped arc of $\mathcal{X}$ is the same as a usual arc of $\mathscr{W}\!(\mathcal{X})$, the motivic measure on warped arcs is inherited (via bootstrapping) from the motivic measure for usual (untwisted) arcs on stacks. Furthermore via this bootstrapping, the change of variables formula in [SU21, SU23b] is automatically upgraded to the warped setting (see Theorem 1.4). In other words, our bootstrapping technique (Theorem 1.1) allows us to sidestep the need to prove a motivic change of variables formula in the warped setting from scratch.
@@ 1.2. Definitions: warped maps
Let $\mathcal{X}$ be a locally finite type Artin stack with affine diagonal over an algebraically closed field $k$ of characteristic 00.
@@
(1) ${\mathcal{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$$\scriptstyle{\pi}$${\mathcal{X}\times_{Y}D\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$$\scriptstyle{p^{\prime}}$${\mathcal{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${D\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${Y}$
@@ 1.3. Statement of results
see Section 5. By Theorem 1.1, we may immediately apply the motivic change of variables formula from [SU21, SU23b] to $\underline{p}$. We now explain why the change of variables formula applied to $\underline{p}$, as opposed to $p$, behaves particularly nicely in light of the non-properness of $\mathcal{X}$.
@@
For the rest of the introduction, we additionally fix the following set-up. Recall from [SU23a, Definition 1.3] that a map $p\colon\mathcal{X}\to Y$ to an algebraic space is called strongly birational if there is a dense open subspace $U\subset Y$ over which $p$ is an isomorphism. We refer to the largest such $U$ as the strong birationality locus.
Notation 1.2.
@@
Using Notation 1.2, we let the canonical locus $\mathcal{C}_{\mathcal{X}}\subset|\mathscr{L}(\mathscr{W}\!(\mathcal{X}))|$ be the set of isomorphism classes of warped arcs $(\pi\colon\mathcal{D}\to D,f\colon\mathcal{D}\to\mathcal{X})$ such that
@@
Using Notation 1.2, for every field $k^{\prime}$, the map
@@ 1.4. Applications to motivic integration
We emphasize again a key conceptual ingredient leading to our motivic change of variables formula Theorem 1.4 (see also Theorem 10.1): since Theorem 1.1 implies that warped maps from $T$ to $\mathcal{X}$ are the same as usual maps $T\to\mathscr{W}\!(\mathcal{X})$, we may immediately bootstrap our motivic change of variables formula in [SU21, SU23b] to warped setting.
@@
Using Notation 1.2, assume $Y$ is irreducible and $U$ is smooth. For any measurable function $f\colon\mathscr{L}(Y)\to\mathbb{Z}\cup\{\infty\}$ with $\mathbb{L}^{f}$ integrable on $\mathscr{L}(Y)$, we have
@@
For simplicity, we stated Theorems 1.4 and 10.1 where $\mathcal{X}\to Y$ is a good moduli space map. However, it applies equally well to any stacky resolution given by the composition of a good moduli space map and a proper map; indeed one simply combines the above theorem with the usual valuative criterion for proper maps. We recall that by [SU23a], if $Y$ has log-terminal singularities, such a resolution can always be chosen to be crepant. Thus any motivic integral over arcs of a log-terminal variety can be expressed canonically as a motivic integral over warped arcs of a crepant stacky resolution of singularities.
We end the introduction by giving a McKay correspondence for linearly reductive groups. Batyrev’s McKay correspondence [Bat99] (conjectured by Reid [Rei92]) is given for quotients $\mathbb{A}^{n}/G$ where $G$ is a finite subgroup of $\operatorname{SL}_{n}$. For algebraic groups $G$, we need an additional constraint which appears as well in the work of Špenko–van den Bergh [SvdB17] on non-commutative crepant resolutions (NCCRs).
To state our McKay correspondence, we must first recall some terminology. If $Y$ is Gorenstein, let $\!\mathscr{J}_{Y}$ be the unique ideal sheaf such that the image of $\Omega_{Y}^{\dim Y}\to\omega_{Y}$ is given by $\!\!\mathscr{J}_{Y}\omega_{Y}$. Denef–Loeser [DL99] introduced the Gorenstein measure
$\mu_{Y}^{\mathrm{Gor}}(C):=\int_{C}\mathbb{L}^{\operatorname{ord}_{\!\mathscr{J}_{Y}}}\mathrm{d}\mu_{Y}$
and proved that the Hodge–Deligne specialization of $\mu_{Y}^{\mathrm{Gor}}(\mathscr{L}(Y))$ recovers the stringy Hodge numbers of $Y$. Recall also the notion of generic $G$-representations.
Definition 1.6 ([SvdB17, Definition 1.6]).
@@
If $V$ is a generic representation of a linearly reductive group $G\subset\operatorname{SL}(V)$, then
$\mu_{V/G}^{\mathrm{Gor}}(\mathscr{L}(V/G))\ =\ \int_{\mathcal{C}_{[V/G]}}\mathbb{L}^{\operatorname{ord}_{\!\mathscr{J}_{V\!/G}}\!\circ\mathscr{L}(\underline{p})-\operatorname{ht}_{\mathscr{W}\!([V/G])/(V/G)}}\mathrm{d}\mu_{\mathscr{W}\!([V/G]),\dim(V/G)}$
where $p\colon[V/G]\to V/G$ denotes the good moduli space map.
@@
When $G$ is a finite subgroup of $\operatorname{SL}(V)$, the genericity assumption on $V$ is automatic.
@@ Acknowledgements.
This work benefited greatly from conversations with many individuals. It is our pleasure to thank Dan Abramovich, Anton Geraschenko, Changho Han, Jesse Kass, Dhruv Ranganathan, Karl Schwede, Wanchun Shen, Karen Smith, Jerry Wang, and Jonathan Wise. We thank Dan Abramovich for sharing with us toric examples that served as initial inspiration for the notion of warped arcs. We thank many colleagues who helped us with naming conventions, especially Dhruv Ranganathan for suggesting that we use the term warp. We are grateful to Dori Bejleri and Dan Edidin for their insights which led to Proposition 4.3. We are indebted to Jack Hall for his patient explanations of many technical details that played a key role in Theorem 1.1(1).
@@ 2. Warped arcs: preliminaries
Throughout this section, we assume the set-up in Notation 1.2 and let $\mathcal{U}:=p^{-1}(U)$. Fix a field extension $k^{\prime}/k$ and let $D=\operatorname{Spec}k^{\prime}[[t]]$. We begin by showing that the canonical arcs introduced in Definition 1.1 are indeed warped arcs.
@@
${\mathcal{D}^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$$\scriptstyle{\pi^{\prime}}$$\scriptstyle{f^{\prime}}$${\mathcal{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{f}$${\mathcal{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${D\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${Y}$
@@
${\mathcal{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{\pi}$${\mathcal{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${D\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Y}$
@@
Let $\mathcal{X}=[\mathbb{A}^{2}/\mathbb{G}_{m}]$ where $\mathbb{G}_{m}$ acts with weights $(1,-1)$. If $\mathbb{A}^{2}$ has coordinates $x,y$, then $\mathcal{X}\to\mathbb{A}^{1}=\operatorname{Spec}k[xy]$ is a good moduli space map. Let $\varphi\colon D=\operatorname{Spec}k^{\prime}[[t]]\to\mathbb{A}^{1}$ be the map sending $xy$ to $t$. Then one computes that $\varphi^{\operatorname{can}}$ is given by the map
@@
where $\mathbb{G}_{m}$ acts on $x,y$ with weights $1,-1$.
@@
The following example played a key role in [SU23a], where we constructed crepant resolutions of all log-terminal singularities using stacks. Similar to Example 2.1, suppose $\mathcal{X}=[\mathbb{A}^{r^{2}}/\operatorname{SL}_{r}]$ where $\mathbb{A}^{r^{2}}$ is viewed as the space of $r\times r$ matrices and $\operatorname{SL}_{r}$ acts by left multiplication. Then the map $\mathcal{X}\to\mathbb{A}^{1}$ sending a matrix to its determinant is a good moduli space map. If $\varphi\colon D=\operatorname{Spec}k^{\prime}[[t]]\to\mathbb{A}^{1}$ is the map sending the coordinate on $\mathbb{A}^{1}$ to $t$, then $\varphi^{\operatorname{can}}$ is given by the map

### v3

# Beyond twisted arcs: a McKay correspondence for reductive groups

@@ 1. Introduction
If $Y$ is a log-terminal variety with quotient singularities, its stringy Hodge numbers have a beautiful interpretation as orbifold Hodge numbers of a smooth Deligne–Mumford stack $\mathcal{X}$ where $\mathcal{X}\to Y$ is a canonical crepant resolution. There are two major obstructions to finding a similar interpretation for the stringy Hodge numbers when $Y$ has worse than quotient singularities. The first is that $Y$ rarely admits a crepant resolution by a Deligne–Mumford stack. In [SU24], we were able to overcome this obstruction by constructing crepant resolutions for all log-terminal varieties using Artin stacks. This leads to the second major obstruction: Artin stacks (other than Deligne–Mumford stacks) are not separated, creating major difficulties in applying previous techniques in motivic integration. In this paper we introduce a versatile framework that overcomes this second obstruction.
@@ 1.1. Overview
While crepant resolutions by varieties or Deligne–Mumford stacks are rare, we proved in recent work [SU24] that crepant resolutions by Artin stacks hold in great generality: every log-terminal variety $Y$ admits a crepant resolution by a smooth Artin stack $\mathcal{X}$. Since log-terminal singularities are precisely the class for which Batyrev defined stringy Hodge numbers [Bat98], this shows that resolutions by smooth Artin stacks provide a natural setting by which to study stringy invariants. Furthermore, log-terminal singularities are a natural class arising in the context of the minimal model program (MMP). Since motivic integration and the MMP are largely controlled by maps from discs and curves, for applications to birational geometry and motivic McKay correspondences, it is particularly important to understand how maps from curves or arcs on $Y$ lift to twisted versions on $\mathcal{X}$. One is therefore in need of new tools for lifting curves and arcs via resolutions by Artin stacks.
@@
Furthermore, we provide a versatile framework for working with warped maps. Indeed, while the valuative criterion and the lack of properness for good moduli spaces provides the impetus for the construction of warped maps, the most prominent feature of our paper is the construction of an Artin stack $\mathscr{W}\!(\mathcal{X})$ which parameterizes warped maps, i.e., a warped map from $T$ to $\mathcal{X}$ is given by a usual map $T\to\mathscr{W}\!(\mathcal{X})$, see Theorem 1.9. Thus, this allows one to bootstrap results from usual maps to warped maps. This bootstrapping technique is essential in our main applications to motivic integration and the McKay correspondence (Theorems 1.18 and 1.21), and is analogous to Olsson’s work in logarithmic geometry [Ols03], where he showed that logarithmic morphisms $S\to Z$ are the same as usual morphisms $S\to\mathcal{L}og_{Z}$ to an auxiliary stack, and hence properties of log maps are inherited from properties of usual maps to stacks.
@@
With the theory of warped maps in hand, we prove a motivic change of variables formula for Artin stacks (see Theorem 1.18) and give a McKay correspondence for linearly reductive groups (see Theorem 1.21). Our change of variables formula for Artin stacks applies, in particular, to crepant resolutions of all log-terminal singularities, as constructed in our previous work [SU24]. While other motivic change of variables formulas for stacks have appeared in the literature, they are all either in the Deligne–Mumford setting or involve making choices, e.g., choosing lifts to arcs of $\mathcal{X}$. Such choices do not always exist and even when they do, there can be infinitely many such choices. On the other hand, our motivic change of variables formula here is completely canonical: by our valuative criterion (Theorem 1.15) outside of a set of measure zero, any arc $\varphi$ of $Y$ lifts canonically to a warped arc $\varphi^{\operatorname{can}}$ of $\mathcal{X}$. Our motivic change of variables formula then expresses any motivic integral over arcs $\varphi$ of $Y$ as an explicit motivic integral over the warped arcs $\varphi^{\operatorname{can}}$ of $\mathcal{X}$.
@@
We wish to emphasize that in addition to our motivic change of variables formula being a novel result, the proof technique is unlike any that previously existed, even when one restricts to the case of Deligne–Mumford stacks. Indeed, for Yasuda to construct a motivic measure for twisted arcs, he needed to first construct spaces parametrizing twisted jets and arcs and then get enough explicit control of these spaces in order to construct the measure. In comparison, our main theorem (Theorem 1.9) showing $\mathscr{W}\!(\mathcal{X})$ is an Artin stack allows us to completely sidestep the need to construct a motivic measure for warped arcs: since a warped arc of $\mathcal{X}$ is the same as a usual arc of $\mathscr{W}\!(\mathcal{X})$, the motivic measure on warped arcs is inherited (via bootstrapping) from the motivic measure for usual (untwisted) arcs on stacks. Furthermore via this bootstrapping, the change of variables formula in [SU21, SU23] is automatically upgraded to the warped setting (see Theorem 1.18). In other words, our bootstrapping technique (Theorem 1.9) allows us to sidestep the need to prove a motivic change of variables formula in the warped setting from scratch.
@@ 1.2. Definitions: warped maps
Let $\mathcal{X}$ be a locally finite type Artin stack with affine diagonal over an algebraically closed field $k$ of characteristic 0.
@@
[…]
@@ 1.3. Statement of results
see Section 5. By Theorem 1.9, we may immediately apply the motivic change of variables formula from [SU21, SU23] to $\underline{p}$. We now explain why the change of variables formula applied to $\underline{p}$, as opposed to $p$, behaves particularly nicely in light of the non-properness of $\mathcal{X}$.
@@
For the rest of the introduction, we additionally fix the following set-up. Recall from [SU24, Definition 1.3] that a map $p\colon\mathcal{X}\to Y$ to an algebraic space is called strongly birational if there is a dense open subspace $U\subset Y$ over which $p$ is an isomorphism. We refer to the largest such $U$ as the strong birationality locus.
Notation 1.13.
@@
Using Notation 1.13, we let the canonical locus $\mathcal{C}_{\mathcal{X}}\subset|\mathscr{L}(\mathscr{W}\!(\mathcal{X}))|$ be the set of isomorphism classes of warped arcs $(\pi\colon\mathcal{D}\to D,f\colon\mathcal{D}\to\mathcal{X})$ such that
@@
Using Notation 1.13, for every field $k^{\prime}$, the map
@@ 1.4. The McKay correspondence
We emphasize again a key conceptual ingredient leading to our motivic change of variables formula Theorem 1.18 (see also Theorem 10.1): since Theorem 1.9 implies that warped maps from $T$ to $\mathcal{X}$ are the same as usual maps $T\to\mathscr{W}\!(\mathcal{X})$, we may immediately bootstrap our motivic change of variables formula in [SU21, SU23] to warped setting.
@@
Using Notation 1.13, assume $Y$ is an irreducible scheme and $U$ is smooth. For any measurable function $f\colon\mathscr{L}(Y)\to\mathbb{Z}\cup\{\infty\}$ with $\mathbb{L}^{f}$ integrable on $\mathscr{L}(Y)$, we have
@@
For simplicity, we stated Theorems 1.18 and 10.1 where $\mathcal{X}\to Y$ is a good moduli space map. However, it applies equally well to any stacky resolution given by the composition of a good moduli space map and a proper map; indeed one simply combines the above theorem with the usual valuative criterion for proper maps. We recall that by [SU24], if $Y$ has log-terminal singularities, such a resolution can always be chosen to be crepant. Thus any motivic integral over arcs of a log-terminal variety can be expressed canonically as a motivic integral over warped arcs of a crepant stacky resolution of singularities.
We end the introduction by giving the McKay correspondence for linearly reductive groups. Batyrev’s McKay correspondence [Bat99] (conjectured by Reid [Rei92]) is given for quotients $\mathbb{A}^{n}/G$ where $G$ is a finite subgroup of $\operatorname{SL}_{n}$. For algebraic groups $G$, we need an additional constraint which appears as well in the work of Špenko–van den Bergh [vVdB17] on non-commutative crepant resolutions (NCCRs).
To state our McKay correspondence, we must first recall some terminology. If $Y$ is log-terminal, let $m$ be such that $Y$ is $m$-Gorenstein and let $\mathscr{J}_{Y,m}$ be the unique ideal sheaf such that the image of $(\Omega_{Y}^{\dim Y})^{\otimes m}\to\omega_{Y,m}$ is given by $\mathscr{J}_{Y,m}\omega_{Y,m}$. Denef–Loeser [DL99] introduced the Gorenstein measure
$\mu_{Y}^{\mathrm{Gor}}(C):=\int_{C}\mathbb{L}^{(1/m)\operatorname{ord}_{\mathscr{J}_{Y,m}}}\mathrm{d}\mu_{Y}\in\widehat{\mathcal{M}}_{k}[\mathbb{L}^{1/m}]$
and proved that the Hodge–Deligne specialization of $\mu_{Y}^{\mathrm{Gor}}(\mathscr{L}(Y))$ recovers the stringy Hodge numbers of $Y$. Recall the notion of generic $G$-representations.
Definition 1.20 ([vVdB17, Definition 1.6]).
@@
If $V$ is a generic representation of a linearly reductive group $G$ such that $(\det V)^{\otimes n}$ is the trivial representation for some $n>0$, then $V/G$ is log-terminal and
$\mu_{V/G}^{\mathrm{Gor}}(\mathscr{L}(V/G))\ =\ \int_{\mathcal{C}_{[V/G]}}\mathbb{L}^{(1/m)\operatorname{ord}_{\!\mathscr{J}_{V\!/G,m}}\!\circ\mathscr{L}(\underline{p})-\operatorname{ht}_{\mathscr{W}\!([V/G])/(V/G)}}\mathrm{d}\mu_{\mathscr{W}\!([V/G]),\dim(V/G)}$
where $p\colon[V/G]\to V/G$ denotes the good moduli space map and where $m$ is such that $V/G$ is $m$-Gorenstein.
@@
The classical McKay correspondence is for finite subgroups of $\operatorname{SL}(V)$. When $G$ is a linearly reductive subgroup scheme of $\operatorname{SL}(V)$, the $G$-representation $\det V$ is trivial (so we may take $n=1$); when $G$ is a finite subgroup of $\operatorname{SL}(V)$, the genericity assumption on $V$ is also automatic.
@@ Acknowledgements.
This work benefited greatly from conversations with many individuals. It is our pleasure to thank Dan Abramovich, Kohei Aoyama, Anton Geraschenko, Changho Han, Jesse Kass, Dhruv Ranganathan, Karl Schwede, Wanchun Shen, Karen Smith, Jerry Wang, and Jonathan Wise. We thank Dan Abramovich for sharing with us toric examples that served as initial inspiration for the notion of warped arcs. We thank many colleagues who helped us with naming conventions, especially Dhruv Ranganathan for suggesting that we use the term warp. We are grateful to Dori Bejleri and Dan Edidin for their insights which led to Proposition 4.4. We are indebted to Jack Hall for his patient explanations of many technical details that played a key role in Theorem 1.9(1).
@@ 2. Warped arcs: preliminaries
Throughout this section, we assume the set-up in Notation 1.13 and let $\mathcal{U}:=p^{-1}(U)$. Fix a field extension $k^{\prime}/k$ and let $D=\operatorname{Spec}k^{\prime}[[t]]$. We begin by showing that the canonical arcs introduced in Definition 1.8 are indeed warped arcs.
@@
[…]
@@
[…]
@@
Let $\mathcal{X}=[\mathbb{A}^{2}/\mathbb{G}_{m}]$ where $\mathbb{G}_{m}$ acts with weights (1,1)(1,-1). If $\mathbb{A}^{2}$ has coordinates $x,y$, then $\mathcal{X}\to\mathbb{A}^{1}=\operatorname{Spec}k[xy]$ is a good moduli space map. Let $\varphi\colon D=\operatorname{Spec}k^{\prime}[[t]]\to\mathbb{A}^{1}$ be the map sending $xy$ to $t$. Then one computes that $\varphi^{\operatorname{can}}$ is given by the map
@@
where $\mathbb{G}_{m}$ acts on $x,y$ with weights 1,11,-1.
@@
The following example played a key role in [SU24], where we constructed crepant resolutions of all log-terminal singularities using stacks. Similar to Example 2.6, suppose $\mathcal{X}=[\mathbb{A}^{r^{2}}/\operatorname{SL}_{r}]$ where $\mathbb{A}^{r^{2}}$ is viewed as the space of $r\times r$ matrices and $\operatorname{SL}_{r}$ acts by left multiplication. Then the map $\mathcal{X}\to\mathbb{A}^{1}$ sending a matrix to its determinant is a good moduli space map. If $\varphi\colon D=\operatorname{Spec}k^{\prime}[[t]]\to\mathbb{A}^{1}$ is the map sending the coordinate on $\mathbb{A}^{1}$ to $t$, then $\varphi^{\operatorname{can}}$ is given by the map